Find the equation of the line which is perpendicular to 3x-7y=-4 and passes through the point (-6,1)
Answer by mananth(16946)

3x-7y=-4 and passes through the point (-6,1)
3 x -7 y = -4
Find the slope of this line

-7 y = -3 x -4
Divide by -7
y = 3/ 7 x + 0.57
Compare this equation with y=mx+b, m= slope & b= y intercept
slope m = 3/7

The slope of a line perpendicular to the above line will be the negative reciprocal -2.33
m1*m2=-1
The slope of the required line will be -7/3

m= -7/3 ,point (-6,1)
Find b by plugging the values of m & the point in
y=mx+b
1 = 14.00 + b
b= -13
m= -7/3
The required equation is y =-7/3 x-13
m.ananth@hotmail.ca